OpenStudy (anonymous):

Solve the system of linear equations below. 2x + 3y = 2 x + 6y = 4

OpenStudy (anonymous):

Solve for a variable in one of the equations, then plug it into the other. well solve for x in the second equation cause its the easiest. x+6y=4 x=-6y+4 so if x is -6y+4 we can plug that into the first equation where x is 2(-6y+4)+3y=2 can you solve for y now?

OpenStudy (anonymous):

y =2/3 ?

OpenStudy (anonymous):

Yep, now plug that into one of the equations for y and solve for x

OpenStudy (anonymous):

would x equal zero?

OpenStudy (anonymous):

or 2

OpenStudy (anonymous):

zero

OpenStudy (anonymous):

thank you :)